### v3

# Inferring Resource-Oriented Intentions using LLMs for Static Resource Leak Detection

Abstract
Resource leaks, caused by resources not being released after acquisition, often lead to performance issues and system crashes. Existing static detection techniques rely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions to find unreleased resources, suffering from both (1) false negatives caused by the incompleteness of predefined resource acquisition/release APIs and (2) false positives caused by the incompleteness of resource reachability validation identification.
To overcome these challenges, we propose InferROI, a novel approach that leverages the exceptional code comprehension capability of large language models (LLMs) to directly infer resource-oriented intentions (acquisition, release, and reachability validation) in code. InferROI first prompts the LLM to infer involved intentions for a given code snippet, and then incorporates a two-stage static analysis approach to check control-flow paths for resource leak detection based on the inferred intentions. We evaluate the effectiveness of InferROI in both resource-oriented intention inference and resource leak detection. Experimental results on the DroidLeaks and JLeaks datasets demonstrate InferROI achieves promising bug detection rate (59.3% and 64.8%) and false alarm rate (18.6% and 24.0%). Compared to three industrial static detectors, InferROI detects 14~45 and 167~503 more bugs in DroidLeaks and JLeaks, respectively. When applied to real-world open-source projects, InferROI identifies 26 unknown resource leak bugs, with 7 of them being confirmed by developers. Finally, manual annotation indicated that InferROI achieved a precision of 74.6% and a recall of 81.8% in intention inference, covering more than 60% resource types involved in the datasets. The results of an ablation study underscores the importance of combining LLM-based inference with static analysis.
I Introduction
Resource leaks, stemming from the failure to release acquired resources (e.g., unclosed file handles), represent critical software defects that can lead to runtime exceptions or program crashes. Such leaks are pervasive in software projects [1] and even appear in online code examples, even for some accepted answers in Stack Overflow posts [2].
To address this issue, researchers have proposed various automated techniques [3, 4, 5, 6, 7] for resource leak detection based on static analysis. These techniques primarily rely on two essential components. The first step involves identifying pairs of resource acquisition APIs and their corresponding resource release APIs; Subsequently, the code is analyzed to verify whether the release API is not appropriately invoked after the acquisition API on the control-flow paths, before the acquired resources become unreachable. For instance, consider the API pair for managing locks: LockManager.acquireLock() represents the resource acquisition API, and LockManager.releaseLock() denotes the corresponding resource release API. A resource leak occurs when releaseLock() is not subsequently called after acquireLock().
However, these existing techniques mainly rely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions to find unreleased resources, leading to the following two limitations. (1) False negatives from the incompleteness of predefined resource acquisition/release APIs: the effectiveness of current resource leak detection techniques [3, 4, 5] hinges on the completeness of predefined API pairs, since they cannot detect resource leaks related to un-predefined resource acquisition/release API pairs. For example, the widely adopted detectors (e.g., SpotBugs, Infer, and CodeInspection) are confined to detecting only a small subset of common resource classes in JDK and Android [8] by default. Thus, the resource leaks involving emerging or less common resources (e.g., AndroidHttpClient in apache httpclient library) cannot be detected by these tools if the specific acquisition/release APIs for AndroidHttpClient are not additionally configured into these tools. (2) False positives from the incompleteness of resource reachability validation identification: the accuracy of reachability analysis for resources along control-flow paths significantly influences the occurrence of false alarms in existing detection methods, since unreachable resources would not cause leaks even without being released after acquisition. Approaches like those introduced by Torlak et al. [3] determine resource reachability by scrutinizing null-checking conditions (e.g., cur!=null) within if-statements. However, this approach falls short for certain resources with alternative means of determining reachability, such as through API calls (e.g., !bank.isDisabled()). Consequently, false alarms are triggered for resources that are actually unreachable along specific paths.
In this paper, we propose InferROI, a novel approach that leverages large language models (LLMs) to boost static resource leak detection. Our intuition is that LLMs have been trained on massive code and documentation corpus and thus have exceptional code comprehension capability [9, 10, 11, 12], suggesting their potential for inferring resource-oriented intentions (i.e., the functionalities of resource acquisition/release and resource reachability validation) in any given code snippet. InferROI consists of two stages, i.e., resource-oriented intention inference and lightweight resource leak detection. (1) For a given code snippet, InferROI first prompts the LLM to directly infer the involved resource acquisition, release, and reachability validation intentions, without relying on any prior knowledge of resource API pairs. The intentions generated by LLMs (which are described in natural language) are then formalized into formal expressions, so as to facilitate the subsequent static analysis stage. (2) InferROI then incorporates a two-stage static analysis approach to check control-flow paths for resource leaks based on the inferred intentions. In particular, InferROI first identifies potential leak-risky paths based on the inferred resource acquisition/release APIs and then prunes leak-risky paths with unreachable resources based on the inferred reachability validation, so as to reduce false alarms of resource leak detection.
While our approach is general and not limited to specific programming languages, we currently implement InferROI for Java given its prevalence. We conduct an extensive evaluation to assess the efficacy of InferROI in detecting resource leaks. Firstly, we apply InferROI to 86 and 784 bugs sourced from the DroidLeaks [8] and JLeaks [13] datasets, respectively, to evaluate its effectiveness in resource leak detection. The results reveal InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% for DroidLeaks, and a bug detection rate of 64.8% with a false alarm rate of 24.0% for JLeaks. Compared to three industrial static detectors, InferROI detects 14~45 and 167~503 more bugs in DroidLeaks and JLeaks, respectively, while maintaining a comparable false alarm rate. Secondly, we employ InferROI for resource detection in real-world open-source projects. Our experiments on 100 methods uncover 26 previously unknown resource leaks, 7 of which are confirmed by developers as of the submission time. Finally, manual annotation for resource-oriented intentions indicated that InferROI achieved a precision of 74.6% and a recall of 81.8% in intention inference, covering 19 (67.9%) and 221 (60.1%) resource types involved in DroidLeaks and JLeaks, respectively. The results of an ablation study underscores the importance of combining LLM-based inference with static analysis.
To summarize, this paper makes the following contributions:
  • A novel LLM-based perspective that infers resource-oriented intentions (resource acquisition, release, and reachability validation) in code instead of mechanically matching predefined APIs. The inferred intentions can be applied to boost static resource leak detection techniques.
  • A lightweight detection approach InferROI that combines the LLM-based intention inference and a static two-stage path analysis to detect resource leaks in code.
  • An extensive evaluation that demonstrates the effectiveness of InferROI in inferring resource-oriented intentions and detecting diverse resource leaks in both existing datasets and real-world open-source projects.
II Motivating Example
This section presents a motivating example to underscore the significance of directly inferring resource-oriented intentions in code through a combination of background knowledge of resource management and code context understanding, rather than relying solely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions. The resource-oriented intentions refer to the functionalities of resource acquisition, release, and reachability validation involved in code.
Figure 1: Comparison between Buggy Version (Left) and Fixed Version (Right) Associated with AndroidHttpClient Leak
Figure 1 illustrates the code diff of a commit [14] of fixing a resource leak associated with AndroidHttpClient, as observed in the DroidLeaks dataset. In the buggy version, the API AndroidHttpClient.newInstance(...) is called to acquire an HTTP client resource at line 166, but the corresponding release API AndroidHttpClient.close() is not called, leading to a resource leak. In the fixed version, a finally structure is introduced at lines 184-187 to ensure the proper release of the acquired HTTP client resource. Through this example, we demonstrate the following key insights:
1. Identifying resource acquisition/release more comprehensively. In the aforementioned example, the presence of a resource leak related to the AndroidHttpClient class goes unnoticed by all the eight subject detectors (i.e., Code Inspection [15], Infer [7], Lint [16], SpotBugs [6], Relda2-FS [17], RElda2-FI [17], Elite [18], and Verifier [19]) evaluated in the DroidLeaks dataset. The reason for this oversight lies in the fact that AndroidHttpClient is not included in the predefined API sets used by these detectors. This limitation highlights a fundamental constraint of the existing detection paradigm, which primarily relies on locating API calls for resource acquisition/release through the resolution of API signatures in code (i.e., type inference) and subsequent matching with predefined API names.
However, by leveraging background knowledge encompassing pertinent concepts such as networking and HTTP client, and diligently analyzing the code contexts, which involve relevant terms like URL and HTTPresponse, it becomes evident that the API call AndroidHttpClient.newInstance(...) is specifically intended to acquire an HTTP client resource. This realization motivates the exploration of inferring resource acquisition and release intentions based on code semantics (e.g., the API calls and the surrounding contexts), enabling more comprehensive identification of resource leaks compared to conventional mechanical API matching.
Furthermore, in addition to AndroidHttpClient, a multitude of APIs are available in various libraries/packages that implement the functionality of HTTP client. For instance, within the selenium library [20], there exist related APIs like HttpClient, JdkHttpClient, NettyClient, TracedHttpClient, NettyDomainSocketClient, PassthroughHttpClient, among others. The eight detectors evaluated in DroidLeaks also fail to detect resource leaks across all of them. This failure can be attributed to the detectors’ limited capability of capturing the underlying common characteristics shared by these APIs, which essentially represent the same resource concept, namely, an HTTP client, and its usage. The presence of this limitation further emphasizes the necessity of incorporating background knowledge and code understanding in resource acquisition and release identification.
2. Identifying resource reachability validation more comprehensively. The validation of resource reachability is a crucial factor in determining the occurrence of resource leaks. In the fixed version depicted in Figure 1, the if-condition at line 185 serves to validate the reachability of the acquired HTTP client resource before its release at line 186. If this if-condition for reachability validation is disregarded, a false alarm will be triggered in the false branch of the if-statement, where the acquired resource is actually unreachable.
Existing detection techniques, such as Torlak et al. [3] and Cherem et al. [21], typically identify reachability validation by matching null-checking conditions. While this approach provides a simple and intuitive means to recognize the reachability validation at line 185, it may overlook certain validation intentions present in other code. For specific resource types, dedicated APIs, such as isClosed() and isDisabled(), are employed to perform reachability validation. In a similar manner to resource acquisition/release identification, a more comprehensive identification of such validation intentions necessitates the incorporation of background knowledge pertaining to common validation verbs, in conjunction with a deep understanding of the code contexts, such as the if-conditions. By considering these elements, resource reachability validation can be assessed more comprehensively, enabling a more accurate detection of resource leaks in code.
III Approach
Based on these insights, we propose InferROI, a novel approach that directly Infers Resource-Oriented Intentions in code to boost static resource leak detection, by leveraging the power of large language models (LLMs). In the approach, LLMs function as extensive knowledge bases with exceptional code understanding capabilities [9, 10, 11, 12], rendering them well-suited for inferring resource-oriented intentions.
Figure 2 represents the approach overview of InferROI. For a given code snippet, InferROI first instantiates the prompt template to instruct the LLM in inferring the involved resource acquisition, release, and reachability validation intentions, which does no require any prior information on resource API pairs. The intentions generated by LLMs (which are described in natural language) are then formalized into formal expressions, so as to facilitate the subsequent static analysis stage. By aggregating these inferred intentions, InferROI then proceeds to apply a two-stage static analysis algorithm to analyze the control-flow paths extracted from the code. The algorithm first identifies potential leak-risky paths based on the inferred resource acquisition/release APIs, and then prunes leak-risky paths with unreachable resources based on the inferred reachability validation across these paths on sibling branches, so as to reduce false alarms of resource leak detection.
Figure 2: Approach Overview of InferROI
III-A Resource-Oriented Intentions
We focus on the following three resource-oriented intentions involved in code, which play a crucial role in resource management.
Resource Acquisition. Statements acquire resources (e.g., locks or database connections), which are normally API invocations and involved with objects/variables of resources (e.g., lock or db). To formally express the intentions of resource acquisition, we introduce the notation $\bm{ACQ(var,lineno)}$, where $var$ denotes the variable storing the acquired resource, and $lineno$ indicates the line number containing the statement responsible for acquiring the resource.
Resource Release. Likewise, statements release previously acquired resources. We formalize the intentions of resource release as $\bm{REL(var,lineno)}$, where $var$ corresponds to the variable storing the resource to be released, and $lineno$ indicates the line number of the statement responsible for performing the release operation.
Resource Reachability Validation. Conditional statements aim to validate whether acquired resources remain reachable within specific branches. To formalize the intentions of resource reachability validation, we employ the notation $\bm{VAL(var,lineno)}$, where $var$ represents the variable storing the resource being validated, and $lineno$ indicates the line number of the statement responsible for performing the validation operation.
Example 3.1. In the method depicted in Figure 1, the intentions of the API calls at line 167, which aims to acquire a HTTP client resource, can be represented as $ACQ({\texttt{\footnotesize client}},167)$. The intention of the API call at line 186, responsible for releasing the acquired HTTP client resource, can be denoted as $REL({\texttt{\footnotesize client}},186)$. The intention of the if statement at line 185, intended to validate the reachability of the acquired HTTP client resource, can be expressed as $VAL({\texttt{\footnotesize client}},185)$.
III-B LLM-Based Intention Inference
To infer resource-oriented intentions present in code, we first prompt the large language model (LLM) with designed templates and then formalize the LLM-generated answers into formal expressions for the subsequent analysis.
III-B1 Feeding LLM with Prompt
For a given code, we construct a prompt using the template depicted in Figure 3. This prompt is then fed into the LLM (i.e., GPT-4 used in this work) for intention inference. The prompt template is designed following the best practices for prompt engineering by OpenAI [22], consisting of three main components: Task Description & Instructions, Output Format Specification, and Code Placeholder.
Figure 3: Prompt Template
Task Description & Instructions component introduces the background (the first sentence) and outlines the instructions (the remaining sentences) to be executed by the LLM. The prompt includes five sequential instructions, with the first two serving as leading instructions to guide the LLM in smoothing the reasoning chain from a basic starting point. In the leading instructions, the LLM is prompted to perform type inference for the involved objects and determine which types represent leakable resources. Based on the analysis, the LLM is then required to infer the three types of resource-oriented intentions by the follow-up three instructions. Throughout the analysis and inference process, the LLM leverages its extensive background knowledge to understand the code contexts and execute the instructions effectively.
Output Format Specification aims to normalize the format of identified intentions, facilitating subsequent parsing of the outputs.
Code Placeholder is where the given code with line numbers is inserted, creating a specific prompt.
Example 3.2. Applying the template, we construct a prompt for the fixed version (right) illustrated in Figure 1. We then employ the GPT-4 model [23] to generate an answer, as shown in Figure 4. Notably, the GPT-4 model exhibits the capability to effectively analyze code semantics and discern resource-oriented intentions. Within the generated answer, resource acquisition intentions are highlighted in red , resource release intentions in green , and resource reachability validation intentions in cyan .
Figure 4: Answer Generated by GPT-4 Model
III-B2 Extracting Formal Expressions
To extract formal expressions of the intentions identified by the LLM, we employ matching patterns that correspond to the output format specification specified in the template. Given an LLM-generated answer, we utilize regular expressions such as “line <line number>: <API call> acquires <resource variable> resource” to extract the line number, API call, and resource variable involved in resource acquisition intentions. Similar regular expressions can be used to extract information related to resource release intentions and reachability validation intentions. We then use the extracted information to instantiate the formal expressions of the three types of resource-oriented intentions.
Example 3.3. From the answer illustrated in Figure 4, the three highlighted lines are matched by the matching patterns. Subsequently, the three formal intention expressions are extracted and instantiated as $ACQ({\texttt{\footnotesize client}},167)$, $REL({\texttt{\footnotesize client}},186)$, and $VAL({\texttt{\footnotesize client}},185)$.
III-C Static Resource Leak Detection
We then propose a lightweight static analysis based on the inferred resource-oriented intentions to detect resource leaks in code.
III-C1 Enumerating Control-Flow Paths
Similar to most existing resource leak detection methods [3, 4, 6, 7], our detection approach is also built upon control-flow analysis. We construct control-flow graphs (CFGs) for the code to detect and extract pruned control-flow paths from these graphs.
Given a method-level code snippet, we begin by parsing it into an abstract syntax tree (AST) and then proceed to construct a CFG by traversing the AST nodes and inserting various control-flow edges between the code statements. The procedure is implemented based on the top of an established CFG construction tool [24]. Within the CFG, certain nodes, such as if-statements and switch-statements, give rise to multiple subsequent control-flow branches. These particular nodes are referred to as branch-nodes within the CFG. Next, we enumerate the paths within the CFG, considering the paths between the entry-node and exit-node. To ensure the efficiency of the enumeration process, we employ two pruning operations.
Loop Structures. Loop statements in the CFG can introduce circular paths, leading to infinite-length paths during the enumeration. To prevent this, we adopt a strategy that involves expanding the true-branch of a loop only once.
Resource-Independent Branches. The presence of branches in the CFG can exponentially increase the number of paths during enumeration, but most of the branches are not related to resource management. To address this, for a branch-node, we further prune its branches that are independent of resource acquisition/release and procedural exit (e.g, return-statements).
Example 3.4. For the buggy version in Figure 1, there is only one control-flow path that sequentially executes from line 160 to line 185. For the fixed version in Figure 1, there are two control-flow paths branched by the if-statement of reachability validation at line 185. The two control-flow paths represented in line numbers are: [160-185, 186, 187-190] and [160-185, 187-190].
III-C2 Detecting Resource Leaks
After enumerating the control-flow paths and inferring the resource-oriented intentions, InferROI proceeds with resource leak detection.
For each concerned resource involved in the inferred intentions, the detection process involves a two-stage path analysis, as illustrated in Algorithm 1. This algorithm takes the concerned resource $res$, all the control-flow paths $\mathcal{P}$, and the resource-oriented intention set $\mathcal{I}$ as inputs. The first stage (lines 1-16) is a single-path analysis that initially identifies leak-risky paths of the concerned resource based on the $ACQ$ and $REL$ intentions. The second stage (lines 17-29) is a cross-path analysis that greedily eliminates the false-alarm-introducing branch-nodes by considering the $VAL$ intentions. After completing both stages, the algorithm (lines 31-36) checks if any path has a corresponding status of $true$. If such a path exists, a resource leak of $res$ is reported.
Algorithm 1 Leak Detection via Two-Stage Path Analysis
1:$res$: concerned resource; $\mathcal{P}$: paths; $\mathcal{I}$: intention set;
2$\triangleright$ Stage 1: Single-Path Analysis
3:for $path\in\mathcal{P}$ do
4:   $rd\_counter\leftarrow 0$;
5:   for $node\in path$ do
6:      $ln\leftarrow$ get line number of $node$;
7:      if $ACQ(res,ln)\in\mathcal{I}$ then
8:         $rd\_counter\leftarrow rd\_counter+1$;
9:      else if $REL(res,ln)\in\mathcal{I}$ then
10:         $rd\_counter\leftarrow rd\_counter-1$;
11:      end if
12:   end for
13:   if $rd\_counter>0$ then
14:      $path.risky\leftarrow$ true;
15:   else
16:      $path.risky\leftarrow$ false;
17:   end if
18:end for
19$\triangleright$ Stage 2: Cross-Path Analysis
20:$br\_nodes\leftarrow$ get all branch-nodes in the paths;
21:sort $br\_nodes$ by their line numbers in a descending order
22:for $br\_node\in br\_nodes$ do
23:   $br\_ln\leftarrow$ get line number of $br\_node$
24:   if not $VAL(res,br\_ln)\in\mathcal{I}$ then
25:      continue;
26:   end if
27:   $G_{prefix}\leftarrow$ group paths containing $br\_node$ by path prefixes;
28:   for $g\in G_{prefix}$ do
29:      $(B_{1},B_{2},...)\leftarrow$ group paths in $g$ by branches of $br\_node$;
30:      $\textsc{Propagate}(B_{1},B_{2},...)$;
31:   end for
32:end for
33$\triangleright$ Reporting Leaks
34:for $path\in\mathcal{P}$ do
35:   if $path.risky$ then
36:      report a leak of $res$;
37:      break;
38:   end if
39:end for
1:procedure Propagate($B_{1},B_{2},...$)
2:   if ($\forall p\in B_{i}\mapsto p.risky$) and ($\nexists p\in\underset{j\neq i}{\cup}B_{j}\mapsto p.risky$then
3:      foreach $p\in B_{i}:p.risky\leftarrow$ false;
4:   end if
5:end procedure
Stage 1: Single-Path Analysis. The algorithm conducts an iterative process, sequentially processing each $path$ in $\mathcal{P}$ and traversing the nodes within $path$ (lines 2-17). During this node traversal, a descriptor counter $rd\_counter$ keeps track of the $ACQ$ and $REL$ intentions involving the resource $res$ within $path$ (lines 3-11). If $rd\_counter$ is greater than 0, it indicates that $path$ is a potential leak-risky path for $res$, and as a result, the risky status of $path$ (i.e., $path.risky$) is initialized as $true$ (lines 12-16).
Stage 2: Cross-Path Analysis. As outlined in Section II, the presence of branches, denoted as unreachable branches, where the resource is inaccessible, gives rise to the possibility of false-positive (FP) leak-risky paths. This, in turn, results in false alarms being reported at line 32. During this phase, we conduct cross-path analysis utilizing the inferred reachability validation intentions to eliminate the false alarms.
Intuitively, addressing this problem involves specifically analyzing each branch-node intended for reachability validation and identifying its unreachable branch. However, the task is complicated by the diverse implementations of validation APIs, as discussed in Section II. Some APIs (e.g., isClosed()) return true to indicate that the resources are reachable, while others (e.g., isActive()) indicate the opposite. To overcome this challenge, our approach utilizes a novel heuristic that collectively analyzes all branches of a branch-node to assess whether it introduces false alarms. If affirmative, such branch-nodes are designated as false-alarm-introducing (FAI) nodes. These FAI nodes are characterized by its one branch being completely leak-risky, while the others are completely not leak-risky. A branch is considered completely leak-risky when all the paths belonging to it exhibit a risky status of $true$. The completely leak-risky branch is the unreachable branch, and the paths within it are FP leak-risky paths for $res$. The insight behind this decision is that if the other branch (the completely not leak-risky branch) is the unreachable one, it suggests the presence of more severe functional bugs rather than resource leaks, such as null pointer errors.
To implement this heuristic, our algorithm initiates by identifying all branch-nodes in the CFG and iterates through them from back to front (lines 18-30). During this process, each branch-node $br\_node$ is examined to determine if it involves a $VAL$ intention. If not, the algorithm skips $br\_node$, recognizing that it does not influence the risky status of the paths (lines 21-24). On the contrary, when $br\_node$ involves a $VAL$ intention, the algorithm conducts two grouping operations (line 24 and line 26, respectively). This results in the division of paths containing $br\_node$ into branches $(B_{1},B_{2},...)$. All these branches share a common path prefix (a subpath from the entry-node to $br\_node$) but belong to different branches of $br\_node$. This implies that, at a specific execution path to $br\_node$, $(B_{1},B_{2},...)$ represent potential subsequent branches of $br\_node$. Upon the branches, the algorithm performs the heuristic by executing the Propagate procedure and updating the risky status of FP leak-risky paths to $false$ (line 28).
Example 3.5. In the buggy version depicted in Figure 1, the detection algorithm successfully identifies a resource leak of the variable client. In contrast, for the fixed version shown in Figure 1, the algorithm identifies a leak-risky path [160-185, 187-190] of the variable client in Stage 1. However, the Update procedure is then executed in Stage 2, which effectively eliminates this path as an FAI path. Consequently, no resource leak of client is reported in this version, indicating that the detection algorithm correctly avoids false alarms and provides a more accurate analysis of resource leaks.
IV Implementation
The current implementation of InferROI utilizes the widely adopted GPT-4 as the LLM and sets the model temperature as 0 to remove generation randomness. During the static resource leak detection, we take into account certain Java-specific language features, such as try-with-resources statements. For the try-with-resources statement, the acquired resources are automatically closed, and there is no need for manual resource closure [25]. To address this, we incorporate a simple rule-based post-processing step to filter out any false alarms introduced by the try-with-resources statements.
V Evaluation
We conduct a comprehensive set of experiments to evaluate the effectiveness of InferROI in detecting resource leaks by employing the DroidLeaks dataset [8] and the JLeaks dataset [13]. InferROI typically takes an average of 3 to 8 seconds to process each code snippet.
First, we apply InferROI to detect resource leaks in the buggy-version and fixed-version code snippets in the datasets, comparing its bug detection rate and false alarm rate against those of baseline detectors (RQ1). By doing so, we assess whether InferROI can surpass the performance of the baseline detectors in accurately identifying resource leaks. In the next part of our evaluation (RQ2), we explore the efficacy of InferROI in detecting resource leaks when applied to real-world open-source projects. This investigation enables us to gauge InferROI’s performance under real-world conditions. Finally, we manually annotate the resource-oriented intentions presented in code to measure the extent to which InferROI successfully infers these intentions (RQ3.a), and conduct an ablation study to analyze and compare the contributions of two key components of InferROI (RQ3.b).
All the research questions are listed as follows.
  • RQ1 (Effectiveness in Leak Detection): How effectively can InferROI detect resource leaks in the DroidLeaks and JLeaks datasets? Can it outperform baseline detectors in this regard?
  • RQ2 (Effectiveness in Project Scanning): How effective is InferROI in detecting previously-unknown resource leak bugs when applied to open-source projects?
  • RQ3.a (Accuracy in Intention Inference): To what degree can InferROI accurately infer the resource-oriented intentions in code?
  • RQ3.b (Ablation Study): What is the contribution of the intention inference and static analysis in InferROI?
V-A Effectiveness in Resource Leak Detection (RQ1)
In this evaluation, we assess the effectiveness of InferROI in detecting resource leaks using two datasets.
V-A1 Data
We utilize two datasets, namely DroidLeaks [8] and JLeaks [13], constructed by analyzing commits dedicated to fixing resource leak bugs. Each bug in the datasets is associated with a concerned resource type, which is determined by reviewing the corresponding bug-fixing commit. By retrieving the bug-fixing commit and the previous commit, both the buggy and fixed versions are obtained for each bug.
  • DroidLeaks: This dataset comprises resource leak bugs sourced from open-source Android applications. Originally, the DroidLeaks dataset included 116 bugs used to assess the performance of eight widely-used resource leak detectors [15, 7, 16, 6, 17, 18, 19]. In our evaluation, we attempt to acquire all 116 bugs, successfully obtaining both the buggy and fixed versions for 86 of them. Consequently, the final collection comprises 172 code snippets covering 28 different resource types, such as Cursor.
  • JLeaks: This is a more recent dataset containing resource leak bugs from open-source Java projects. The original JLeaks dataset included 1,094 bugs used to evaluate the performance of three widely-used resource leak detectors [6, 7, 26]. For our evaluation, we filtered out instances based on two conditions: 1) the resource type is annotated, and 2) both the buggy and fixed versions are free of syntax errors. This filtering process resulted in 784 bugs (1,568 code snippets) encompassing 368 different resource types.
V-A2 Baselines
We compare InferROI with three well-maintained industrial static detectors, as evaluated in previous studies [8, 13].
  • SpotBugs [6]: SpotBugs, formerly known as FindBugs, is a static analysis tool that identifies bugs and potential issues in Java programs.
  • Infer [7]: Infer is a static analysis tool developed by Facebook that detects various types of programming errors in C, C++, and Java code.
  • PMD [26]: PMD is a widely-used static code analysis tool that identifies potential issues in code.
V-A3 Procedure
Following previous study [8], we adopt the metrics Bug Detection Rate and False Alarm Rate to evaluate the detection effectiveness of InferROI and baseline detectors.
TABLE I: Resource Leak Detection on DroidLeaks Dataset
Detector # Detected Bugs (Bug Detection Rate) # False Alarms (False Alarm Rate)
SpotBugs 6 (6.9%) 0 (0%)
Infer 37 (43.0%) 16 (18.6%)
PMD 10 (11.6%) 6 (7.0%)
InferROI (ours) 51 (59.3%) 16 (18.6%)
TABLE II: Resource Leak Detection on JLeaks Dataset
Detector # Detected Bugs (Bug Detection Rate)
SpotBugs* 44 (24.3%)
Infer* 5 (9.4%)
PMD 341 (43.5%)
InferROI (ours) 508 (64.8%)
    Infer and SpotBugs encounter compilation issues with the majority of test instances.
For each bug in our experiments, we apply InferROI to its buggy version and fixed version. If InferROI reports a resource leak in the buggy version related to the corresponding concerned resource type, it is considered to have correctly detected the bug. Conversely, if InferROI reports a resource leak in the fixed version for the corresponding concerned resource type, it signifies a false alarm. The two metrics are calculated as follows:
Bug Detection Rate=# detected bugs in buggy versions# total experimented $\textit{Bug Detection Rate}=\frac{\text{\# detected bugs in buggy versions}}{\text{\# total experimented bugs}}$
False Alarm Rate=# reported false alarms in fixed versions# total experimented $\textit{False Alarm Rate}=\frac{\text{\# reported false alarms in fixed versions}}{\text{\# total experimented bugs}}$
V-A4 Results
The evaluation results for DroidLeaks and JLeaks are presented in Table I and Table II, respectively.
DroidLeaks: InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% across 86 bugs in DroidLeaks. Compared to the baseline detectors, InferROI achieves notably higher bug detection, detecting 14~45 more bugs, while maintaining a comparable false alarm rate.
JLeaks: InferROI showcases a bug detection rate of 64.8% with a false alarm rate of 24.0% across 784 bugs in JLeaks. In contrast to the baseline detectors, InferROI achieves significantly superior bug detection, identifying 167~503 more bugs in JLeaks. It’s important to note that we do not report the false alarm rate of the baselines due to the unavailability of data necessary for its calculation in current JLeaks dataset.
Analysis: (i) Bug Detection Analysis: The higher bug detection rate of InferROI can be attributed to the advantages of the resource-oriented intention inference utilized by InferROI, as discussed in the insights provided in Section II (see RQ3.a for the resource coverage assessment). However, it is essential to acknowledge that the resource-oriented intention inference cannot entirely replace the value of more sound program analysis techniques. For instance, in the case of buggy versions not detected by InferROI, a substantial portion is caused by the intricacies of Android’s lifecycle management and callback mechanisms. This observation indicates an important future direction, i.e., to explore better integration of LLM-based resource-intention inference with more advanced program analysis techniques, which could enable more general and powerful resource leak detection. (ii) False Alarm Analysis: False alarms reported by InferROI primarily stem from incorrect or missing intentions. When InferROI incorrectly identifies an acquisition intention or overlooks a release intention, a false alarm occurs. To address this issue, we intend to explore the integration of LLM-based prompt engineering techniques, such as in-context learning and chain-of-thought [27], to enhance the capabilities of LLMs. Notably, a significant portion of these “false alarms” arises from misclassifications due to the quality of the ground truth. Since both datasets are curated based on bug-fixing commits, certain “fixed” versions still contain resource leaks. This underscores the ongoing challenge of ensuring the quality of resource leak datasets.
Summary: InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% across 86 bugs in DroidLeaks, and a bug detection rate of 64.8% with a false alarm rate of 24.0% across 784 bugs in JLeaks. This performance shows a promising potential of detecting resource leaks based on resource-oriented intentions.
V-B Effectiveness in Open-Source Project Scanning (RQ2)
We apply InferROI on real-world Java projects to evaluate its efficacy in detecting previously-unknown resource leaks.
V-B1 Data
We crawl 115 Java open-source projects with more than 50 stars, which are created after December 31, 2021. The date filtering condition is used to avoid the projects are included in the training data of the employed GPT-4, which is trained using a huge amount of data up to September 2021. To save time and the cost of calling the GPT-4 API (https://platform.openai.com/docs/api-reference), we randomly filter 100 methods from 13 projects, by matching 20 common resource terms reported in [28], instead of completely scaning these projects. The employed terms are listed as follows: stream, reader, client, writer, lock, player, connection, monitor, gzip, ftp, semaphore, mutex, stream, camera, jar, buffer, latch, socket, database, scanner, cursor.
V-B2 Procedure
We apply InferROI on these 100 methods to detect resource leaks. For each reported leak, we first manually annotate whether it is a true bug by reading code and consulting related information. For the true bugs, we then submit the corresponding bug-fix pull requests and help the project developers review them.
V-B3 Results
In the 100 methods, InferROI reports 29 resource leaks and 26 are annotated as true bugs. At the submission time, 7 of the true bugs have been confirmed by the project developers, and the pull requests have been merged.
Comparison with Static Detector: The static detector, PMD, reports 20 resource leaks in the 100 methods, with 18 of them being true bugs. Among these 18 bugs identified by PMD, InferROI successfully detects 17 of them. The lone bug missed by InferROI involves a method comprising over 400 lines of code, posing challenges for LLM processing. Conversely, InferROI identifies 9 additional true bugs not caught by PMD. Noteworthy among these are resource types such as ManagedBuffer from the Apache Spark library, InMemoryDirectoryServer from UnboundID LDAP, and a custom type JDBCConnection specific to the project.
Comparison with Resource Mining Approach: Of the 26 bugs, 3 are beyond the scope of the state-of-the-art resource acquisition/release API pair (RAR pair) mining approach MiROK [28]. MiROK primarily concentrates on extracting the definitions of abstract RAR pairs, which differs from our objective of directly detecting resource-oriented intentions in the usages of RAR pairs. Nonetheless, the abstract RAR pairs mined by MiROK can aid in filtering suspicious methods for InferROI, similar to the 20 resource terms employed in this evaluation.
Summary: InferROI successfully detects 26 previously-unknown resource leak bugs in 100 methods from real-world open-source projects. The results demonstrate the practical capability of InferROI in effectively detecting resource leaks in real-world software projects.
V-C Accuracy in Resource-Oriented Intention Inference (RQ3.a)
In this evaluation, we evaluate the effectiveness of InferROI’s LLM-based resource-oriented intention inference.
V-C1 Procedure
To evaluate the effectiveness of InferROI’s intention inference, two authors independently annotate all three types of resource-oriented intentions present in the 172 code snippets collected from the DroidLeaks dataset. In cases of disagreement, they engage in discussions to reach a final decision, ensuring consistent and accurate annotations. These manual annotations serve as the ground truth intentions ($GTs$) for the evaluation.
Subsequently, we apply InferROI to the code snippets to infer the intentions (denoted as $Preds$), and then calculate the precision and recall of the inference based on the manual annotations. Precision represents the proportion of intentions inferred by InferROI that align with the ground truth, while recall indicates the proportion of true intentions from the ground truth that InferROI correctly infers. The precision and recall calculations are as follows:
$\textit{precision}=\frac{|GTs\cap Preds|}{|Preds|};\textit{recall}=\frac{|GTs\cap Preds|}{|GTs|}$
Furthermore, we evaluate the resource coverage of InferROI for DroidLeaks and JLeaks and compare it with PMD. For InferROI, we determine the portion of the 28 (DroidLeaks) and 368 (JLeaks) resource types covered by the inferred intentions.
V-C2 Results
InferROI achieves a precision of 74.6% and a recall of 81.8% in resource-oriented intention inference, indicating its ability to relatively accurately and comprehensively identify intentions within the code snippets. Moreover, InferROI demonstrates superior versatility in addressing a broader spectrum of resource types, attaining a coverage of 67.9% (19 types) for DroidLeaks and 60.1% (221 types) for JLeaks, compared to PMD’s coverage of 32.1% (9 types) and 12.8% (47 types), respectively.
Through analysis of the evaluation results, we attribute the effectiveness of InferROI’s intention inference to the powerful capabilities of LLMs in leveraging background knowledge and understanding code context. For example, InferROI can identify the AndroidHttpClient-related intentions, which are ignored by other baseline detectors.
Summary: InferROI exhibits a precision of 74.6% and a recall of 81.8% in inferring resource-oriented intentions, achieving resource type coverages of 67.9% and 60.1% for DroidLeaks and JLeaks, respectively. This achievement can be attributed to the capabilities of LLMs, coupled with the efficacy of our designed prompt template.
V-D Ablation Study (RQ3.b)
We conduct an ablation study to analyze and compare the contributions of two key components of InferROI: the LLM-based resource-oriented intention inference and the static resource leak detection.
V-D1 Procedure
To perform the ablation study, we create two prompt templates to enable GPT-4 to directly detect resource leaks in code, resulting in two resource leak detectors: GPTLeak and GPTLeak-chain. The prompt templates for GPTLeak and GPTLeak-chain are illustrated in Figure 5, and they are structured similarly to the template utilized by InferROI. For GPTLeak, the instruction is straightforward, while for GPTLeak-chain, an additional detection instruction (the last sentence) is incorporated into the template used by InferROI to provide specific guidance to the GPT-4 model during the decision-making process. We apply both GPTLeak and GPTLeak-chain detectors to the 172 code snippets collected from the DroidLeaks dataset. Subsequently, we calculate the bug detection rate and false alarm rate for each detector.
Figure 5: Prmopt Templates for the Two GPT-based Detectors
V-D2 Results
As shown in Table III, InferROI substantially outperforms both GPTLeak and GPTLeak-chain, achieving the highest bug detection rate and the lowest false alarm rate. These findings show the effectiveness of InferROI’s combination of the LLM-based intention inference and the static analysis in detecting resource leaks. While LLMs offer extensive knowledge and understanding of code, they may still lack robust reasoning capacity to accurately detect resource leaks. On the other hand, the static analysis also provides a more rigorous approach, which can address the limitations of LLMs and reduce the potential for false positives and false negatives in the detection process. By synergistically combining the strengths of both approaches, InferROI achieves a balanced and effective resource leak detection mechanism, ultimately enhancing the overall accuracy and reliability of its results.
An interesting observation is that GPTLeak-chain performs worse than GPTLeak when utilizing the instruction chain for the intention inference. This phenomenon may suggest that the LLMs detect resource leaks using different logic than the procedure of static analysis-based detection. Explicitly inferring the three types of resource-oriented intentions proves to be beneficial for boosting static resource leak detection, but appears to be unnecessary for LLMs. This observation emphasizes the significance of prompt design when applying LLMs for resource leak detection.
TABLE III: Ablation Study Results
Detector # Detected Bugs (Bug Detection Rate) # False Alarms (False Alarm Rate)
GPTLeak 44 (51.2%) 18 (20.9%)
GPTLeak-chain 32 (37.2%) 36 (41.9%)
InferROI (ours) 51 (59.3%) 16 (18.6%)
Summary: The results of the ablation study confirm the contributions of both LLM-based resource-oriented intention inference and static resource leak detection in InferROI.
VI Discussion
VI-A Effects of Different LLMs
In our current implementation, InferROI utilizes the powerful LLM, GPT-4, which has demonstrated its effectiveness in inferring resource-oriented intentions based on the designed prompt template. It is important to consider the potential impact of using different LLMs with varying model parameter sizes and training dataset scales. For instance, LLMs such as GPT-3.5 might exhibit different performance levels with GPT-4. To gain insights into this aspect, we conducted some preliminary experiments with GPT-3.5 to qualitatively assess its effectiveness in resource-oriented intention inference. Our observations indicate that GPT-3.5 exhibits a notable performance gap, with approximately 20% lower precision and recall compared to GPT-4. This difference could be attributed to the more powerful analytical capabilities of GPT-4 in code understanding, as highlighted in recent investigation [9, 10]. These findings motivate further exploration of developing specific LLMs tailored for resource-oriented intention inference, potentially involving fine-tuning of “smaller” models like Llama [29, 30]. Moreover, although we set the temperature of the LLM as zero to remove randomness, the LLM outputs are still sometimes unstable. Therefore, we further repeat our experiments on a small set of cases and find the variance in the final detection performance is negligible.
VI-B Static Analysis
InferROI adopts a lightweight intraprocedural static analysis for detecting resource leaks. Instead of utilizing interprocedural analysis, we opt for an intraprocedural analysis based on the observation that many lower-level resource APIs (e.g., AndroidHttpClient) are encapsulated within higher-level resource APIs (e.g., Socket). By directly inferring intentions for higher-level resources, we can reduce the complexity of interprocedural program analysis during the detection process.
Nevertheless, it is important to acknowledge the limitations in the employed static analysis, especially in scenarios where certain types of resource leaks might remain undetected. For instance, resource leaks occurring within Android callbacks represent a challenging area where our lightweight static analysis may fall short. Our key insights lie in directly inferring resource-oriented intentions based on code context understanding to boost static leak detection. The advantage of inferring resource-oriented intentions is its potential applicability to various static leak detection techniques, including callback-aware analyses like Relda2 [17]. By integrating the inferred resource-oriented intentions into such existing techniques, we can enhance their capability to handle more comprehensive cases. In this regard, our LLM-based resource-oriented intention inference can provide complementary insights and improvement for existing static resource leak detection, with the potential to empower a wide range of analysis methods to achieve more comprehensive and accurate results.
VI-C Threats to Validity
The internal validity of our studies is potentially affected by the randomness in data sampling and the subjectiveness in data annotation. To mitigate these threats, we adopted commonly-used data sampling strategies and involved multiple annotators to minimize any preference bias. Another concern is the risk of data leakage, where the evaluation data might have been incorporated into the training data of the utilized LLM. However, the ablation study (RQ3.b) already shows that purely querying GPT-4 itself to detect resource leak has poor effectiveness, indicating that the effectiveness of InferROI is mainly credited to our proposed intention inference and static analysis algorithm not to the data leakage. Furthermore, we further curate projects created after the date of the GPT-4’s training data collection (in RQ2) to mitigate this issue. As for the external validity, it is essential to consider the evaluation data used in our work, which might not guarantee the full generality of our findings. To address this concern, we evaluate InferROI on both existing datasets and real-world open-source projects to detect resource leaks.
VII Related Work
Resource Leak Detection. Automated resource leak detection techniques [3, 4, 5, 8] have been proposed to detect whether some resource is not being released after its acquisition. Typically there are two important components for resource leak detection. First, identify the potential API pairs for resource acquisition and resource release; then based on the RAR pairs, analyze the code to check whether the release API is not subsequently called after the acquisition API. The majority of existing resource leak detection techniques are concentrated on the analysis part by proposing more precise and more scalable code analysis approaches [3, 4, 5, 31, 32, 33]. For example, Torlak et al. [3] combine intra-procedural analysis and inter-procedural analysis to enable more scalable and more accurate detection for system resource leaks (e.g., I/O stream and database connections); Wu et al. [4] propose an inter-procedural and callback-aware static analysis approach to detect resource leak in Android apps; Kellogg et al. [5] incorporate ownership transfer analysis, resource alias analysis, and fresh obligation creation to enable more precise analysis. The API pairs used in most existing techniques [3, 4] are often predefined by human expertise and heuristic rules, which not only require non-trivial human efforts but also have limited coverage in libraries and APIs. For example, Torlak et al. [3] manually collect API pairs and detect resource leaks for stream and database in JDK. In addition, SpotBugs [6] only considers the predefined stream related API pairs and thus could only detect stream related resource leaks. More recently, Bian et al. [34] propose SinkFinder, which mines and classifies frequent resource-related API pairs for a given project.
Different from the existing detection approaches, our work does not rely on prepared API pairs and directly infers resource-oriented intention in code by utilizing the exceptional code comprehension capability of large language models.
LLMs for Software Engineering. Recent studies have demonstrated that pre-trained language models can be utilized as knowledge bases by using synthetic tasks similar to the pre-training objective to retrieve the knowledge/information stored in the models [35, 36]. These works have shown that language models can recall factual knowledge without any fine-tuning by using proper prompts. Several works have focused on exploring effective prompt templates to improve performance of PLMs on downstream problems. In a recent survey by Liu et al. [37], prompt templates are broadly classified into two categories: cloze prompts [38, 35], which entail filling in the blanks in text or code, and prefix prompts [39, 40], which continue generating content following a specified prefix. Recent studies have explored the application of prompt learning in various software engineering tasks. Wang et al. [41] investigated the effectiveness of prompt learning in code intelligence tasks such as clone detection and code summarization. Huang et al. [42] used prompt learning for type inference.
In our work, we consider LLMs as knowledge bases that contain resource management knowledge with powerful code understanding capability, and prompt LLMs to infer resource-oriented intentions for resource leak detection.
VIII Conclusions and Future Work
In this work, we propose InferROI, a novel resource leak detection approach which boosts static analysis via LLM-based resource-oriented intention inference. Given a code snippet, InferROI prompts the LLM in inferring involved intentions from the code. By aggregating these inferred intentions, InferROI utilizes a lightweight static-analysis based algorithm to analyze control-flow paths extracted from the code, thereby detecting resource leaks. We evaluate InferROI for the Java programming language and investigate its effectiveness in resource leak detection. Experimental results on the DroidLeaks and JLeaks datasets demonstrate that InferROI exhibits promising bug detection rate (about 60%) and false alarm rate (about 20%). Moreover, when applied to open-source projects, InferROI identifies 26 unknown resource leak bugs. In the future, we will explore the extension of InferROI to other programming languages and integrate InferROI with more program analysis techniques.